OpenStudy (anonymous):

verify:2cos^2(x/2)=sin^2x/1-cosx

OpenStudy (anonymous):

I think you can start it with RHS: \[\frac{\sin^2(x)}{1 - \cos(x)}\] As \(sin^2(x) + cos^2(x) = 1\) \[\frac{1 - \cos^2(x)}{1 - \cos(x)}\] Now just factorize numerator by using: \[\frac{(1-\cos(x))(1+\cos(x))}{(1 - \cos(x))}\] Just cancel whatever is being cancelled and then use: \[1 + \cos(x) = 2 \cos^2(\frac{x}{2})\]

OpenStudy (anonymous):

I think in LHS, there must be 2cos^2(x/2), check that..

OpenStudy (anonymous):

oh yep your right i accidently left that part out im glad you caught it